what is the energy stored in a capacitor

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
The energy stored in a capacitor can be calculated from the equivalent expressions:
U=\frac{1}{2}\frac{Q^{2}}{C}=\frac{1}{2}QV=\frac{1}{2}CV^{2}
where
Q= charge
V= potential
C= capacitance
This energy is stored in the form of electric field.